OpenStudy (anonymous):

Solve the following system of equations. 7x+2y=5 2x+3y+16

myininaya (myininaya):

There is a couple of ways to go about this Which way(s) do you know?

myininaya (myininaya):

we can do elimination or substitution?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

7x+2y=5 ---- (1) 2x+3y+16=0 2x+3y=-16 ----(2) now u understand?

myininaya (myininaya):

Oh wait you are missing an equal sign in the second *equation*

OpenStudy (anonymous):

oh you're right!

OpenStudy (anonymous):

7x+2y=5 2x+3y=16

myininaya (myininaya):

So I was saying thre is two ways And I was asking what way of the two ways I mention do you prefer? (Note I do not recommend graphing to solve a system of liner equations because sometimes it ends up being an approximation)

myininaya (myininaya):

I think aamal was talking about the elimination way :)

myininaya (myininaya):

I'm correct to say she? We want to add the equations together such that one of the variables zeros out when doing so Do you know an equation still holds if you multiply both sides by the same thing?

myininaya (myininaya):

Like I like to work with the small numbers Look at the coefficients of y in both equations Can you tell me what they are?

OpenStudy (anonymous):

2 and 3?

myininaya (myininaya):

yes exactly What do you get when you multiply those numbers?

OpenStudy (anonymous):

6

myininaya (myininaya):

So is there a number we can multiply the first equation by such that the coefficient of y in the first equation is 6?

myininaya (myininaya):

You told me 2(3)=6 right?

OpenStudy (anonymous):

yes i did

myininaya (myininaya):

ok great! So if we multiply the first equation by 3 we will get the coefficient of y in the first equation to be 6 like so 21x+6y=15 Do you see how everything in the first equation got multiplied by 3?

OpenStudy (anonymous):

yes i do see that

myininaya (myininaya):

Do you know what we will multiply the second equation by? Remember we are trying to get both y's to have coefficient 6.

OpenStudy (anonymous):

2?

myininaya (myininaya):

yes! :) so what would the second equation look like?

OpenStudy (anonymous):

\[2(2x+3y=16)\] is that right?!

myininaya (myininaya):

yes

myininaya (myininaya):

So.... that would be what after distributing the two?

OpenStudy (anonymous):

i honestly don't know

myininaya (myininaya):

Ok so when I multiplied everything in the first equation by 3 this is what I did: 3(7x+2y=5) 3(7x+2y)=3(5) 3(7x)+3(2y)=3(5) 3(7)x+3(2)y=3(5) 21x+6y=15 I multiplied each term by 3 -------- So you wrote that we need to do 2(2x+3y=16) 2(2x+3y)=2(16) 2(2x)+2(3y)=2(16) Can you finish this part?

OpenStudy (anonymous):

\[4x+6y=16\]

myininaya (myininaya):

well almost 2(16)=32 right?

OpenStudy (anonymous):

oh yes!

myininaya (myininaya):

So we have the two equations rewritten as: 21x+6y=15 4x+6y=32 right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

Now all you do is take the opposite of one of these equations (it doesn't matter which but just do one) What I mean by opposite is that we will just change the sign in front of each term... This will elimination the y terms when we add the equations together

myininaya (myininaya):

Do you know what I mean?

OpenStudy (anonymous):

yes i know what you mean so it will now look like \[-4x-6y=-32\] am i right?

myininaya (myininaya):

So the opposite of (4x+6y=32) is -4x-6y=-32 And like I said just do this for one of the equations (NOT BOTH and it doesn't matter which you choose) omg perfect! that is the exact equation I used lol

myininaya (myininaya):

Now lets add these monsters!

OpenStudy (anonymous):

wait why would you add?!

myininaya (myininaya):

So first let me just write what we have 21x+6y=15 -4x-6y=-32 ------------- When you add these it will eliminate the y term Like we will add the x terms 21x-4x=17x We will add the y terms 6y-6y=0y=0 We will add the constants (that is the number that is not being multiplied by a variable) 15-32=-17 --------- So we have 21x+6y=15 -4x-6y=-32 ------------- 17x =-17 Do you see what I mean?

OpenStudy (anonymous):

oh yeah i see now!

myininaya (myininaya):

So now we are trying to solve 17x=-17 Do you know how to solve this?

OpenStudy (anonymous):

divide by 17 on both sides

myininaya (myininaya):

Yes great! :)

OpenStudy (anonymous):

which =-1

myininaya (myininaya):

yes x is -1 ! :)

OpenStudy (anonymous):

so x=-1 ?!

OpenStudy (anonymous):

yayy! thank you so much!

myininaya (myininaya):

Now .... we are almost done....

myininaya (myininaya):

We need to solve for y

myininaya (myininaya):

Can you do that? Since we know x is -1 then we should be able to replace x with -1 in either equation which will mean we will only have one unknown in that equation which is a very good thing....

myininaya (myininaya):

So we had 7x+2y=5 as one of the equations ... We know x is -1 so we will replace x with -1 like so 7(-1)+2y=5

myininaya (myininaya):

You could have chose either equation you started out with, it doesn't matter. Just use one though. All you need is one. You can do both but that is extra work.

myininaya (myininaya):

Do you see that we only have one unknown in the equation 7(-1)+2y=5?

OpenStudy (anonymous):

yes i do

myininaya (myininaya):

Ok do you know how to solve this for y?

OpenStudy (anonymous):

no i don't know

myininaya (myininaya):

Go ahead and apply order of operations to the left hand side of the equal sign Like multiplication comes first out of the operations given in the left hand side So we do 7 times -1 which is -7 So we have -7+2y=5 right? Maybe this looks better to you...?

OpenStudy (anonymous):

yeah i did that!

myininaya (myininaya):

Now you can't add -7 and 2y because -7 and 2y are not liker terms so you have done all you could to that one side

myininaya (myininaya):

Great! :)

myininaya (myininaya):

So we have -7+2y=5 And we want to isolate the y So we must figure out some way to undo the operations around the y to get it by itself Now remember whatever you do to one side of the equation you must do to the other

myininaya (myininaya):

So I usually like to undo additions and subtractions before undoing the multiplication and divisions

OpenStudy (anonymous):

add the -7 to 5

myininaya (myininaya):

Well add 7 on both sides (not -7) Do you know what we want to add 7 on both sides and not -7?

myininaya (myininaya):

why*

OpenStudy (anonymous):

wouldn't y=6?!

myininaya (myininaya):

yep :) You did it! Go you! :)

OpenStudy (anonymous):

thanks so much! you really helped me to understand and especially because i have my algebra exams on thursday i know im going to need to know this!

myininaya (myininaya):

Good luck! Practice practice! :)

OpenStudy (anonymous):

i will! and thanks :)